OpenStudy (anonymous):

(y)=2(x+4) what are the points?

OpenStudy (anonymous):

This is a line.. so infinite points can lie over it.. but as a sample these two points may be taken (0,8) and (1, 10)

OpenStudy (anonymous):

would the points be (4,8)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

put x=4, and y=8 r u getting LHS = RHS..?? if no then points do not lie

OpenStudy (anonymous):

i dont know what LHS+RHS is? and oh it would be (8,4)

OpenStudy (anonymous):

=**

OpenStudy (anonymous):

LHS id left hand side and RHS is right hand side of "=" sign in the equation

OpenStudy (anonymous):

i dont understand any of this....

OpenStudy (anonymous):

would they be (2,4)?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

put any value of X and simplify for Y

OpenStudy (anonymous):

I DONT KNOW HOW TO DO THAT! :/

OpenStudy (anonymous):

(y)=2(x+4) putting x=2, y= 2(2+4) or y= 2(6) or y= 12 so for x=2, y=12 hence the point can be (2, 12)